# Causal Effect Estimation under Networked Interference without Networked Unconfoundedness Assumption

### v1

@@ Abstract
Estimating causal effects under networked interference is a crucial yet challenging problem. Existing methods based on observational data mainly rely on the networked unconfoundedness assumption, which guarantees the identification of networked effects. However, the networked unconfoundedness assumption is usually violated due to the latent confounders in observational data, hindering the identification of networked effects. Interestingly, in such networked settings, interactions between units provide valuable information for recovering latent confounders. In this paper, we identify three types of latent confounders in networked inference that hinder identification: those affecting only the individual, those affecting only neighbors, and those influencing both. Specifically, we devise a networked effect estimator based on identifiable representation learning techniques. Theoretically, we establish the identifiability of all latent confounders, and leveraging the identified latent confounders, we provide the networked effect identification result. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
Machine Learning, ICML
1 Introduction
Estimating causal effects under network interference is a crucial yet challenging problem across various domains, including human ecology (Ferraro et al., 2019), advertising (Parshakov et al., 2020), and epidemiology (Barkley et al., 2020). The key challenge is that networked interference introduces interactions between units, violating the Stable Unit Treatment Value Assumption (SUTVA). For example, when evaluating the effect of a flu vaccine on individual infection rates, a standard causal inference approach assumes that an individual’s infection risk depends only on their own vaccination status. However, in reality, vaccination generates herd immunity effects—vaccinating one person may reduce disease transmission within the population, indirectly lowering the infection risk of others. This violation of SUTVA introduces bias into traditional causal inference methods, rendering standard estimands inapplicable (Forastiere et al., 2021). To model the interference between units, the existing methods focus on estimating three kinds of networked effects: main effects (effects of units’ own treatments), spillover effects (effects of units’ treatments on other units), and total effects (combined main and spillover effects).
@@ 1 Introduction
To estimate causal effects from observational networked data, a series of works have been proposed under the networked unconfoundedness assumption. This assumption posits that no unobserved confounders exist beyond the observed covariates and the covariates of neighboring units. Under this assumption, Forastiere et al. (2021) establish the identification of network effects and propose the joint generalized propensity score for effect estimation. Building on this, Chin (2019); Ma & Tresp (2021); Cai et al. (2023) introduce balanced representation techniques to construct conditional outcome estimators for effect estimation. Additionally, Liu et al. (2019); Chen et al. (2024) develop doubly robust estimators to improve the robustness of network effect estimation under network interference.
@@
Figure 2: Assumed causal graph in this paper. $x$ denotes observed proxies, $u$ denotes latent confounders, $t$ denotes the treatment, and $y$ denote the outcome of interest. We assume latent confounders $u$ contains three types of variables, i.e., $u^{i}$ affecting unit itself, $u^{n}$ affecting unit’s neighborhoods, and $u^{c}$ affecting both.
@@
However, the networked unconfoundedness assumption is often violated in real-world scenarios, significantly limiting the effectiveness of existing methods. For example, in the case of flu vaccination, whether a person chooses to get vaccinated may depend on their income level or their family’s financial situation. However, such socioeconomic factors are often difficult to measure directly due to privacy concerns or data collection limitations. In such cases, latent confounders exist, violating the networked unconfoundedness assumption and introducing bias into existing methods.
@@
To tackle the above challenge, we aim to develop a method that does not rely on the networked unconfoundedness assumption. Specifically, we begin by exploring three types of latent confounders, shown in Figure 2, which hinder the effect identification: $u^{i}$ affecting only the individual, $u^{n}$ affecting only neighbors, and $u^{c}$ influencing both. Rather than assuming networked unconfoundedness, we investigate the identifiability of latent confounders in the presence of networked interference. We found that networked interference provides additional auxiliary information that facilitates the identification of latent confounders. Built on the identified latent confounders, we theoretically establish the networked effect identification result and further devise an effect estimator under networked interference. Overall, our contribution can be summarized as follows:
@@
  • We address the problem of networked effect identification and estimation in the presence of latent confounders. We categorize three types of latent confounders that hinder identification.
  • We explore the identifiability of latent confounders under networked interference, leveraging which, we achieve the networked effect identification.
@@
  • We devise an estimator built on the theoretical findings. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
2 Related Works
@@
Classic Causal Inference has been studied in two languages: the graphical models (Pearl, 2009) and the potential outcome framework (Rubin, 1974). The most related method is the propensity score method in the potential outcome framework, e.g., IPW method (Rosenbaum & Rubin, 1983; Rosenbaum, 1987), which is widely applied to many scenarios (Rosenbaum & Rubin, 1985; Li et al., 2018; Cai et al., 2024). There are also many outcome regression models, including meta-learners (Künzel et al., 2019), neural networks-based works (Johansson et al., 2016; Assaad et al., 2021). By incorporating them, one can construct a doubly robust estimator (Robins et al., 1994), i.e., the effect estimator is consistent as either the propensity model or the outcome repression model is consistent.
@@
Causal Inference without SUTVA has drawn increasing attention recently. Liu et al. (2016) extend the traditional propensity score to account for neighbors’ treatments and features and propose a generalized Inverse Probability Weighting (IPW) estimator. Forastiere et al. (2021) define the joint propensity score and then propose a subclassification-based method. Drawing upon previous works, Lee et al. (2021) consider two IPW estimators and derive a closed-form estimator for the asymptotic variance. Based on the representation learning, Ma & Tresp (2021) add neighborhood exposure and neighbors’ features as additional input variables and applies HSIC to learn balanced representations. Jiang & Sun (2022) use adversarial learning to learn balanced representations for better effect estimation. Ma et al. (2022) propose a framework to learn causal effects on a hypergraph. (Cai et al., 2023) propose a reweighted representation learning method to learn balanced representations. Under networked interference, McNealis et al. (2023); Liu et al. (2023); Chen et al. (2024) propose an estimator to achieve DR property. However, these works do assume the unconfoundedness assumption, which might not hold in real-world scenarios. Different from them, we explore the problem of networked effect estimation without the unconfoundedness assumption.
@@
Causal Inference without Uncounfoundedness Assumption is an important problem since the unconfoundedness assumption is usually violated in observational studies. Classic methods to solve this problem usually assume there exist additional variables, e.g., instrumental variable (Pearl et al., 2000; Stock & Trebbi, 2003; Wu et al., 2022), proximal variable (Miao et al., 2018; Tchetgen Tchetgen et al., 2024). Another effective way to address this problem is to recover the latent confounder using representation learning methods. CEVAE (Louizos et al., 2017) assumes that latent confounders can be recovered by their proxies and applies VAE to learn confounders. As a follow-up work, TEDVAE (Zhang et al., 2021) decouples the learned latent confounders into several factors to achieve a more accurate estimation of treatment effects. In the mediation analysis, DMAVAE (Xu et al., 2023) proposes to recover latent confounders using the VAE similar to CEVAE. Our work is closely related to these works. Different from them, we focus on the causal effect without the unconfoundedness assumption in the presence of networked interference. We also provide theoretical guarantees for the latent confounder identifiability, which ensures the effectiveness of our estimator.
3 Notations, Assumptions, Esitimands
@@
In this section, we start with the notations used in this work. Let $U\in\mathcal{U}$ be the latent confounders and also let $X\in\mathcal{X}$ be the proxies of latent confounders. Let $T\in\{0,1\}$ denote a binary treatment, where $T=1$ indicates a unit receives the treatment (treated) and $T=0$ indicates a unit receives no treatment (control). Let $Y\in\mathcal{Y}$ be the outcome. We assume that $U$ can be decomposed to $U^{i}$ affecting the unit itself, $U^{c}$ affecting the neighborhood, and $U^{n}$ affecting both. Let lowercase letters (e.g., $x,y,t$) denote the value of random variables. Let lowercase letters with subscript $i$ denote the value of the specified $i$-th unit. Thus, a network dataset is denoted as $D=(\{x_{i},t_{i},y_{i}\}_{i=1}^{n},E)$, where $E$ denotes the adjacency matrix of network and $n$ is the total number of units. We denote the set of first-order neighbors of $i$ as $\mathcal{N}_{i}$ and denote the treatment and feature vectors received by unit $i$’s neighbors as $t_{\mathcal{N}_{i}}$ and $x_{\mathcal{N}_{i}}$. Due to the presence of networked interference, a unit’s potential outcome is influenced not only by its treatment but also by its neighbors’ treatments, and thus the potential outcome is denoted by $y_{i}(t_{i},t_{\mathcal{N}_{i}})$. The observed outcome $y_{i}$ is known as the factual outcome, and the remaining potential outcomes are known as counterfactual outcomes.
@@
Further, following Forastiere et al. (2021); Chen et al. (2024), we assume that the dependence between the potential outcome and the neighbors’ treatments is through a specified summary function $g$: $\{0,1\}^{|\mathcal{N}_{i}|}\rightarrow[0,1]$, and let $z_{i}$ be the neighborhood exposure given by the summary function, i.e., $z_{i}=g(t_{\mathcal{N}_{i}})$. We aggregate the information of the neighbors’ treatments to obtain the neighborhood exposure by $z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}t_{j}}{|\mathcal{N}_{i}|}$. Therefore, the potential outcome $y_{i}(t_{i},t_{\mathcal{N}_{i}})$ can be denoted as $y_{i}(t_{i},z_{i})$, which means that under networked interference, each unit is affected by two kinds of treatments: the binary individual treatment $t_{i}$ and the continuous neighborhood exposure $z_{i}$.
@@
$\psi(t,z):=\mathbb{E}[Y(t,z)],$ (1)
$\mu(t,z,x,x_{\mathcal{N}}):=\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}}],$
Based on the average dose-response function, existing works mostly focus on the following causal effects:
Definition 3.1 (Average Main Effects (AME) ).
@@
Definition 3.2 (Average Spillover Effects (ASE) ).
@@
Definition 3.3 (Average Total Effects (ATE) ).
@@
Definition 3.4 (Individual Main Effects (IME) ).
@@
Definition 3.5 (Individual Spillover Effects (ISE) ).
@@
Definition 3.6 (Individual Total Effects (ITE) ).
@@
Assumption 3.7 (Network Consistency).
@@
Assumption 3.8 (Network Overlap).
@@
Assumption 3.9 (Neighborhood Interference).
@@
These assumptions are commonly assumed in existing causal inference methods such as Forastiere et al. (2021); Cai et al. (2023); Ma et al. (2022). Specifically, Assumption 3.7 states that there can not be multiple versions of a treatment. Assumption 3.8 requires that the treatment assignment is nondeterministic. Assumption 3.9 rules out the dependence of the outcome of unit $i$, $y_{i}$, from the treatment received by units outside its neighborhood, i.e., $t_{j},j\notin\mathcal{N}_{i}$, but allows $y_{i}$ to depend on the treatment received by his neighbors, i.e., $t_{k},k\in\mathcal{N}_{i}$. Also, Assumption 3.9 states the interaction dependence is assumed to be through a summary function $g$. Note that Assumption 3.9 is reasonable in reality for some reason. First, in many applications units are affected by their first-order neighbors, and the affection of higher-order neighbors is also transported through the first-order neighbors. Second, it is also reasonable that a unit is affected by a specific function of other units’ treatment, e.g., how much job-seeking pressure a unit has will depend on how many of its friends receive job training.
@@
Existing methods additionally assume the following assumption:
Assumption 3.10 (Networked Unconfoundedness).
@@
The individual treatment and neighborhood exposure are independent of the potential outcome given the individual and neighbors’ features, i.e., $\forall t,z,\quad y_{i}(t,z){\perp\!\!\!\perp}t_{i},z_{i}|x_{i},x_{\mathcal{N}_{i}}$.
@@
Assumption 3.10 is an extension of the traditional unconfoundedness assumption and indicates that there is no unmeasured confounder which is the common cause of $y_{i}$ and $t_{i},z_{i}$.
@@
Under the assumptions above, the networked effects can be identified (Forastiere et al., 2021; Cai et al., 2023). However, Assumption 3.10 might be too strong to hold, since we can not promise that all confounders are observed in real-world scenarios. Instead, we assume a much weaker assumption by incorporating the latent variables as follows:
Assumption 3.11 (Latent Networked Unconfoundedness).
@@
The individual treatment and neighborhood exposure are independent of the potential outcome given the latent individual and neighbors’ confounders, i.e., $\forall t,z,\quad y_{i}(t,z){\perp\!\!\!\perp}t_{i},z_{i}|u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}$.
@@
This assumption allows for the latent confounders $u^{i},u^{c},u^{n}$. Here, we recognize three types of latent confounders. What serves as an adjustment set under the networked setting is the units’ $u^{i},u^{c}$ and the neighbors’ $u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}$. This also motivates us to identify each latent confounder for a better estimation. In the next section, we will introduce the identifiability of each latent confounder, and further achieve networked effect identification.
4 Networked Causal Effect Identification via Representation Learning

### v2

@@ Abstract
Estimating causal effects under networked interference from observational data is a crucial yet challenging problem. Most existing methods mainly rely on the networked unconfoundedness assumption, which guarantees the identification of networked effects. However, this assumption is often violated due to the latent confounders inherent in observational data, thereby hindering the identification of networked effects. To address this issue, we leverage the rich interaction patterns between units in networks, which provide valuable information for recovering these latent confounders. Building on this insight, we develop a confounder recovery framework that explicitly characterizes three categories of latent confounders in networked settings: those affecting only the unit, those affecting only the unit’s neighbors, and those influencing both. Based on this framework, we design a networked effect estimator using identifiable representation learning techniques. From a theoretical standpoint, we prove the identifiability of all three types of latent confounders and, by leveraging the recovered confounders, establish a formal identification result for networked effects. Extensive experiments validate our theoretical findings and demonstrate the effectiveness of the proposed method.
Index Terms:
Causal effects, networked interference, latent confounders, identifiability.
I Introduction
Estimating causal effects in the presence of networked interference is a crucial yet challenging problem across various domains, including human ecology [1], advertising [2], and epidemiology [3]. The fundamental difficulty lies in that interference between units inherently violates the Stable Unit Treatment Value Assumption (SUTVA), which underpins classical causal inference. For example, when evaluating the effect of a flu vaccine on individual infection rates, a standard causal inference approach assumes that an individual’s infection risk depends only on their own vaccination status. In reality, vaccination induces herd immunity: vaccinating one person may reduce the infection risks of others, leading to indirect or spillover effects (Figure 1). Such a violation of SUTVA introduces bias into traditional causal inference methods, rendering standard estimands inapplicable [4]. To model the interference between units, prior studies focus on estimating three kinds of networked effects: main effects (effects of units’ own treatments), spillover effects (effects of units’ treatments on other units), and total effects (combined main and spillover effects).
@@ I Introduction
To estimate causal effects from observational networked data, a series of methods have been developed under the networked unconfoundedness assumption. This assumption posits that no latent confounders exist beyond the observed covariates and the covariates of neighboring units. Under this assumption, [4] proposes the joint generalized propensity score for effect estimation and establishes the identification of network effects. Subsequent works introduce balanced representation techniques to construct conditional outcome estimators for effect estimation [5, 6, 7] and develop doubly robust estimators to improve the robustness of network effect estimation [8, 9] under networked interference.
@@
Figure 2: Assumed causal graph in this paper. $x$ denotes observed proxies, $u$ denotes latent confounders, $t$ denotes the treatment, and $y$ denote the outcome of interest. We assume that latent confounders $u$ contain three types of variables, i.e., $u^{i}$ affecting the unit itself, $u^{n}$ affecting the unit’s neighbors, and $u^{c}$ affecting both.
@@
However, the networked unconfoundedness assumption is often violated in real-world scenarios, significantly limiting the effectiveness of existing methods. For example, in the case of flu vaccination, whether a person chooses to get vaccinated may depend on their income level or their family’s financial situation. However, such socioeconomic factors are often difficult to measure directly due to privacy concerns or data collection limitations. Such latent confounders, violating the networked unconfoundedness assumption, bias both main and spillover effect estimation, as well as total effect estimation, fundamentally limiting existing approaches.
@@
To address this limitation, we propose a framework that eliminates the need for the networked unconfoundedness assumption. Specifically, we begin by exploring three categories of latent confounders, illustrated in Figure 2, that hinder the effect identification: $u^{i}$ affecting only the unit, $u^{n}$ affecting only neighbors, and $u^{c}$ influencing both. Rather than assuming networked unconfoundedness, we investigate the identifiability of latent confounders in the presence of networked interference. A key insight of our work is that networked interference itself provides auxiliary information that can be exploited to recover latent confounders Building on this recovery, we theoretically establish the networked effect identification result and further devise a practical effect estimator under networked interference. Overall, our contribution can be summarized as follows:
@@
  • We address the problem of networked effect identification and estimation in the presence of latent confounders, and categorize three types of latent confounders that hinder identification.
  • We establish the identifiability of latent confounders by leveraging the auxiliary information revealed by networked interference, and prove that networked effects can be identified.
@@
  • We devise a networked effect estimator built upon these theoretical findings. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
II Related Works
@@
Classic Causal Inference has been studied in two languages: the graphical models [10] and the potential outcome framework [11]. The most related method is the propensity score method in the potential outcome framework, e.g., IPW method [12, 13], which is widely applied to many scenarios [14, 15, 16]. There are also many outcome regression models, including meta-learners [17], neural networks-based works [18, 19, 20, 21]. By incorporating them, one can construct a doubly robust estimator [22], i.e., the effect estimator is consistent as either the propensity model or the outcome repression model is consistent.
@@
Causal Inference without SUTVA has drawn increasing attention recently. [23] extends the traditional propensity score to account for neighbors’ treatments and features and proposes a generalized Inverse Probability Weighting (IPW) estimator. [4] defines the joint propensity score and then proposes a subclassification-based method. Drawing upon previous works, [24] considers two IPW estimators and derives a closed-form estimator for the asymptotic variance. Based on the representation learning, [6] adds neighborhood exposure and neighbors’ features as additional input variables and applies HSIC to learn balanced representations. [25] uses adversarial learning to learn balanced representations for better effect estimation. [26] proposes a framework to learn causal effects on a hypergraph. [7] proposes a reweighted representation learning method to learn balanced representations. Under networked interference, [27, 28, 9] propose an estimator to achieve DR property. Additionally, [29] proposes SPNet to estimate networked effects without the networked unconfoundedness assumption. However, Most works do assume the unconfoundedness assumption, which might not hold in real-world scenarios. Different from them, we explore the problem of networked effect estimation without the unconfoundedness assumption. While [30] also does not assume this assumption, its estimand only focuses on the effects of the unit’s treatment and and it does not consider the identifiability guarantee of latent confounders. Different from [29], we focus on three kinds of networked effects and establish the latent confounder identifiability results, enabling the effect identification.
@@
Causal Inference without Uncounfoundedness Assumption is an important problem since the unconfoundedness assumption is usually violated in observational studies. Classic methods to solve this problem usually assume there exist additional variables, e.g., instrumental variable [31, 32, 33], proximal variable [34, 35]. Another effective way to address this problem is to recover the latent confounder using representation learning methods. CEVAE [36] assumes that latent confounders can be recovered by their proxies and applies VAE to learn confounders. As a follow-up work, TEDVAE [37] decouples the learned latent confounders into several factors to achieve a more accurate estimation of treatment effects. In the mediation analysis, DMAVAE [38] proposes to recover latent confounders using the VAE similar to CEVAE. Our work is closely related to these works. Different from them, we focus on the causal effect without the unconfoundedness assumption in the presence of networked interference. We also provide theoretical guarantees for the latent confounder identifiability, which ensures the effectiveness of our estimator.
III Notations, Assumptions, Estimands
@@
In this section, we start with the notations used in this work. Let $U\in\mathcal{U}$ be the latent confounders and also let $X\in\mathcal{X}$ be the proxies of latent confounders. Let $T\in\{0,1\}$ denote a binary treatment, where $T=1$ indicates a unit receives the treatment (treated) and $T=0$ indicates a unit receives no treatment (control). Let $Y\in\mathcal{Y}$ be the outcome. We assume that $U$ can be decomposed to $U^{i}$ affecting the unit itself, $U^{c}$ affecting the neighbors, and $U^{n}$ affecting both. Let lowercase letters (e.g., $x,y,t$) denote the value of random variables. Let lowercase letters with subscript $i$ denote the value of the specified $i$-th unit. Thus, a network dataset is denoted as $D=(\{x_{i},t_{i},y_{i}\}_{i=1}^{n},E)$, where $E$ denotes the adjacency matrix of network and $n$ is the total number of units. We denote the set of first-order neighbors of $i$ as $\mathcal{N}_{i}$ and denote the treatment and feature vectors received by unit $i$’s neighbors as $t_{\mathcal{N}_{i}}$ and $x_{\mathcal{N}_{i}}$. Due to the presence of networked interference, a unit’s potential outcome is influenced not only by its treatment but also by its neighbors’ treatments, and thus the potential outcome is denoted by $y_{i}(t_{i},t_{\mathcal{N}_{i}})$. The observed outcome $y_{i}$ is known as the factual outcome, and the remaining potential outcomes are known as counterfactual outcomes.
@@
Further, following [4, 9], we assume that the dependence between the potential outcome and the neighbors’ treatments is through a specified summary function $g$: $\{0,1\}^{|\mathcal{N}_{i}|}\rightarrow[0,1]$, and let $z_{i}$ be the neighborhood exposure given by the summary function, i.e., $z_{i}=g(t_{\mathcal{N}_{i}})$. We aggregate the information of the neighbors’ treatments to obtain the neighborhood exposure by $z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}t_{j}}{|\mathcal{N}_{i}|}$. Therefore, the potential outcome $y_{i}(t_{i},t_{\mathcal{N}_{i}})$ can be denoted as $y_{i}(t_{i},z_{i})$, which means that under networked interference, each unit is affected by two kinds of treatments: the binary individual treatment $t_{i}$ and the continuous neighborhood exposure $z_{i}$.
@@
$\psi(t,z)=\mathbb{E}[Y(t,z)],$ (1)
$\mu(t,z,x,x_{\mathcal{N}})=\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}}].$
Based on the above (conditional) average dose-response function, existing works mostly focus on the following causal effects:
Definition 1 (Average Main Effects (AME) ).
@@
Definition 2 (Average Spillover Effects (ASE) ).
@@
Definition 3 (Average Total Effects (ATE) ).
@@
Definition 4 (Individual Main Effects (IME) ).
@@
Definition 5 (Individual Spillover Effects (ISE) ).
@@
Definition 6 (Individual Total Effects (ITE) ).
@@
Assumption 1 (Networked Consistency).
@@
Assumption 2 (Networked Overlap).
@@
Assumption 3 (Neighborhood Interference).
@@
These assumptions are commonly assumed in existing causal inference methods such as [4, 7, 26]. Specifically, Assumption 1 states that there can not be multiple versions of a treatment. Assumption 2 requires that the treatment assignment is nondeterministic. Assumption 3 rules out the dependence of the outcome of unit $i$, $y_{i}$, from the treatment received by units outside its neighbors, i.e., $t_{j},j\notin\mathcal{N}_{i}$, but allows $y_{i}$ to depend on the treatment received by its neighbors, i.e., $t_{k},k\in\mathcal{N}_{i}$. Also, Assumption 3 states the interaction dependence is assumed to be through a summary function $g$. Note that Assumption 3 is reasonable in reality for some reason. First, in many applications, units are affected by their first-order neighbors, and the affection of higher-order neighbors is also transported through the first-order neighbors. Second, it is also reasonable that a unit is affected by a specific function of other units’ treatment, e.g., how much job-seeking pressure a unit has will depend on how many of its friends receive job training.
@@
Existing methods additionally make the following assumption to ensure effect identification:
Assumption 4 (Networked Unconfoundedness).
@@
The individual treatment and neighborhood exposure are independent of the potential outcome given the unit and neighbors’ features, i.e., $\forall t,z$, we have $y_{i}(t,z){\perp\!\!\!\perp}t_{i},z_{i}|x_{i},x_{\mathcal{N}_{i}}$.
@@
Assumption 4 is an extension of the traditional unconfoundedness assumption and indicates that there is no latent confounder that is the common cause between $y_{i}$ and $t_{i},z_{i}$.
@@
Under the assumptions above, the networked effects can be identified [4, 7]. However, Assumption 4 might be too strong to hold, since we can not promise that all confounders are observed in real-world scenarios. Instead, we make a much weaker assumption by incorporating the latent variables as follows:
Assumption 5 (Latent Networked Unconfoundedness).
@@
The individual treatment and neighborhood exposure are independent of the potential outcome given the latent unit’s and neighbors’ confounders, i.e., $\forall t,z$, we have $y_{i}(t,z){\perp\!\!\!\perp}t_{i},z_{i}|u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}$.
@@
This assumption allows for the latent confounders $u^{i},u^{c},u^{n}$. Here, we recognize three categories of latent confounders: $u^{i}_{i}$ affecting the unit itself, $u_{i}^{n}$ affecting the unit’s neighbors, and $u^{c}_{i}$ affecting both. Thus, what serves as a confounder set under the networked setting is the units’ $u^{i},u^{c}$ and the neighbors’ $u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}$. This also motivates us to identify each latent confounder for a better estimation. In the next section, we will introduce the identifiability of each latent confounder and further achieve networked effect identification.
IV Latent Confounder Identifiability Enables Networked Effect Identification
In this section, our goal is to establish networked effect identification in the presence of latent confounders $u^{i},u^{c},u^{n}$. According to Assumption 5, such identification requires access to the latent confounders. This motivates us to employ identifiable causal representation learning techniques for their recovery. To this end, we first introduce the latent confounder identifiability (Theorem 1) and further show that the recovered latent confounders can be disentangled (Theorem 2). Building upon these results, we ultimately establish the identification of networked effects (Theorem 3).